LEP EXPERIMENTS COMMITTEE

Minutes of the 39th meeting,
9 December 1994






OPEN SESSION
 

1. LEP machine status

A. Faugier reported on the very successful LEP operations during 1994 with the 8-bunch pretzel optics. New records were achieved for integrated luminosities
(65 pb-1 per year and 950 nb-1 per day), peak luminosity (2.4 x 1031 cm-2s-1) and overall machine efficiency (60%). The background problems in L3 were eliminated after the addition of extra beam-pipe shielding in September.
 

2. High luminosity in LEP 1 and LEP 2

S. Myers reported the results of the recent bunch-trains tests, and presented the prospects for increased luminosities at both LEP 1 and LEP 2. Collisions of single bunch-trains (1 x 4 bunches) in ALEPH and DELPHI gave luminosities of 5-6 x 1030 cm-2s-1, indicating expected starting performances in 1995
(with 4 x 4 bunches) close to the best previously obtained with pretzel optics. The detectors operated successfully with bunch-trains, and backgrounds were acceptable using asymmetric beam-separation bumps. Whereas LEP 1 operates at the beam-beam limit (current per bunch), it is expected that LEP 2 will be total-current limited (RF-power limited). In consequence the optimum bunch-train configuration at LEP 2 is probably 4 x 2 (or perhaps 4 x 3). The luminosity targets for 1995 are an integrated luminosity of 90 pb-1 per year and a peak luminosity of 3.5 x 1031 cm-2s-1. For LEP 2, the target remains as 500 pb-1
over 3 years, implying a peak luminosity of 7 x 1031 cm-2s-1.
 

3. Transverse polarization in collision at LEP

M. Placidi presented the results of yet another recent success at LEP, performed in the context of the proposal to test longitudinal spin polarization
(LEPC 94-8/P8). A high transverse polarization of the beam (40%) has been maintained during intense collisions (beam-beam tune shift ~0.04) at an energy of peak-2 GeV. Quantitatively, the depolarization time constant was measured to be at least a factor of 3 larger than the polarization time constant.
 

4. Report of the Energy Working Group

R. Jacobsen described the remarkable work of the Energy Working Group in measuring and monitoring the absolute energy of LEP. Effects on the beam energy are now being understood and monitored at the sub-MeV level. The tide of the earthÝs crust (10 MeV peak shift) is one of the largest effects; others range from the water level of Lac Leman (10 MeV) to the daily cycle in the earthÝs magnetic field due to the solar wind (1 MeV). Key requirements during a scan of the Z0 peak are to record as much machine information as possible, to keep the machine stable, and to average over variations. With present knowledge, the overall machine-related errors (i.e. common to all experiments) during a scan are expected to be s (mZ) ~ s (GZ) = 1.0 MeV.
 

5. Status of the LEP 2 superconducting rf system

D. Boussard presented a progress report on the work of the SL-RF group in preparing the LEP 2 sc rf cavities. Niobium coating on copper is still the critical process in the fabrication of cavities. Cavity production is now proceeding smoothly, in spite of its sensitivity to small parameter changes, following improvements in equipment and procedures by the manufacturers. The design of the power couplers has been substantially improved, achieving a factor 6 reduction in internal power dissipation. The designs of the higher-order-mode couplers are also close to completion. Two sc rf modules (a module contains 4 cavities, each of 4 cells) have been successfully operated in LEP at gradients up to the design value of 6 MV/m, for 470 and 770 hours, respectively. In summary, with the detailed designs almost finalized, the sc rf team is now in good shape to study the problem of installing and operating a large sc rf system in LEP.
 

6. Reports on the LEP experiments

L3 :            S. Gentile
OPAL:        R.D. Heuer
DELPHI:    J.E. Augustin
ALEPH:      L. Rolandi
All detectors reported successful operation during 1994. The total statistics are now around 3.5 Z0 hadronic decays (130 pb-1) per experiment. OPAL has almost completed the repair of its Si micro-vertex detector, which had been damaged in an accident in October. All detectors indicated their readiness to operate with bunch-trains at the beginning of 1995 and, in particular, ALEPH indicated progress towards operating with a train length of 740 ns (i.e. operating
efficiently with 4 x 4 bunches). Regarding the requests for LEP running in 1995, all detectors are interested in increased Z0 statistics with bunch-trains
operation and in a scan of the Z0 peak (with an off-peak luminosity of about 20 pb-1). Three detectors requested high-energy running for physics
(a few pb-1 at ~140 GeV) after the October technical stop; L3 places a higher priority in 1995 on increased Z0 statistics, and proposed that high-energy
running for physics be postponed until the start of the 1996 cycle.
 

CLOSED SESSION
 

Present: J. Allaby, W. Bartel, M. Calvetti, T. Camporesi, G. Coignet, J.D. Dowell (Chairman), L. Foà, B. Foster, P. Hansen, W. Hollik,
K. Hübner, P.G. Innocenti, G. Kantardjian, J. Kirkby (Secretary), K.H. Kissler, M. Klein, B. Mansoulié, G. Matthiae, D.J. Miller,
S. Myers, J. Panman, D. Schlatter, D.O. Williams and F. Zwirner.

Apologies: A. Weidberg.

Part Time: A. Blondel, M. Placidi.
 

1. Approval of the minutes of the 38th meeting

The minutes of the 38th meeting (LEPC 94-14, LEPC 38) were approved without modification.
 

2. Chairman's report

In the open session, the Chairman had announced that the first General Meeting of the workshop on LEP 2 physics is scheduled for 2-3 February 1995, in the Main Auditorium at CERN. He had also announced that an audio-visual broadcast of the open session was taking place in real-time on Internet, on an experimental basis, for the benefit of LEP physicists and others who could not attend the session in person. This was the first such broadcast of a CERN committee session (further information can be obtained from Christian Isnard/CN).

In the closed session, the Chairman joined with the committee in congratulating the LEP team for a remarkably successful year: record luminosities, successful bunch-trains tests, observation of transverse polarization in collision, precise calibration and monitoring of the beam energy and, of particular note, great progress on the sc rf system for LEP 2.

The Chairman has received a memorandum from the spokesman of L3 concerning a proposal to construct a very small angle tagger (VSAT) for two-photon physics. The proponents have indicated that a detailed document will be submitted shortly to the LEPC and have requested an early decision in order to allow installation to take place in the 95-96 shutdown. A referee
(M. Klein) was appointed.

Finally, the Chairman reported that the Research Board had approved all the recommendations made at the last LEPC meeting.
 

3. Discussion on bunch-trains

With the positive results of the bunch-trains tests, and with the readiness of the detectors, the committee recommended approval for LEP operation with bunch-trains, commencing at the start of the 1995 cycle. It was noted that the pretzel separators will now remain in place until June, providing an effective safety net in case unforeseen difficulties are encountered with bunch-trains. The committee looks forward with interest to hearing the results of bunch-trains operation of LEP at its June 1995 meeting.
 

4. Discussion of the energy calibration and run plan for 1995

The committee was impressed by the fine work of the Energy Calibration Group. The group is making a vital contribution to LEP 1 physics, and it will have an equally important rôle at LEP 2. The committee encourages the continued active participation of experts from the machine and from each of the four LEP detectors in the Energy Calibration Group.

Regarding the run plan for 1995, the committee recommended a scan of the Z0 peak to take place after the bunch-trains are commissioned. The LEP co-ordinator was asked to co-ordinate the requests of the four detectors and to submit a detailed plan at the June 1995 LEPC meeting. The plan should ensure compatibility between the developments foreseen for the machine and the scanning requirements of a stable, well-monitored machine. Although one detector (L3) requested that high-energy running for physics be postponed until 1996, the committee recommended a high-energy run for the period following the 2nd technical stop in October, assuming the presence of sufficient sc rf cavities. It is hoped that up to 5 pb-1 per experiment could be generated at an energy of about 140 GeV, which would allow a sensitive search for charginos in an unexplored mass range.
 

5. Discussion of the LEP 2 superconducting rf system

The committee congratulated the sc rf team for their substantial progress in 1994. The designs of the main couplers and higher-order mode couplers are almost finalized, cavity manufacture is now reliable, and operational experience of sc cavities in LEP has begun. A first planning for the installation of the Phase II sc cavities will be discussed at the LEP Performance Workshop, Chamonix, January 1995.

S. Myers said that the peak beam energies and luminosities at LEP are under study by the LEP 2 Working Group. At LEP 2 (Phase II), with 192 sc cavities and 56 (i.e. half of the total) copper cavities , a luminosity of about 7 x 1031 cm-2s-1 can be reached at a beam energy of 90 GeV. At LEP 2 (Phase III) - using all 224 sc rf cavities that have been ordered from industry, together with half the copper cavities - this luminosity can be reached at the maximum energy of about 93 GeV.

Since industrial manufacture of cavities for LEP 2 will end in the 2nd half of 1995, the committee asked for a discussion to take place at its June 1995 meeting on the physics and machine aspects related to the peak energy of LEP 2. The primary physics concern is the overlap of the accessible mass ranges for a Higgs between LEP 2 and the LHC. F. Zwirner was appointed to co-ordinate the physics aspects and S. Myers to co-ordinate the machine aspects for this discussion.
 

6. Discussion of the Proposal LEPC 94-8/P8: A Test of Longitudinal Spin Polarization at LEP

The committee congratulated the proponents on their successful demonstration of high (40%) transverse polarization while beams are colliding, which indicates the potential exists for a physics programme at LEP based on longitudinal beam polarization. The referees (W. Bartel, W. Hollik and G. Matthiae) presented further reports, and the proponents (A. Blondel and M. Placidi) were invited to part of the discussion. The importance of the ultimate physics goal - namely the measurement of sin2qW to a precision of 1 x 10-4 from the left-right polarization asymmetry ALR - was underscored once more by the committee. After a long discussion, however, and in view of the conflict of the proposed schedule with the higher priority LEP 2 programme, the committee recommended not to perform the spin rotator test in 1996, as presented in this proposal. Nevertheless the committee encouraged the continuation of tests involving transverse polarization, which are of interest both to this proposal and to the energy calibration of LEP. It was noted that results would be obtained from the spin rotators that are planned for installation in H1 and ZEUS at HERA in 1997. The committee expressed its interest that longitudinal beam polarization remain open as a possible future option beyond LEP 2 and recommended that the matter be reviewed again in 1997 taking into account the planning for the LHC.
 

7. Discussion of the 1995 LEPC meeting at Cogne

The committee discussed the draft agenda for the LEPC meeting at Cogne, 17 - 23 September 1995, to review LEP physics and plans.
 

8. Report from the LEP Co-ordinator

T. Camporesi reported that the 1995 LEP schedule (attached) had been approved by the Research Board.
 

9. A.O.B.

In closing the meeting, the Chairman warmly thanked Jim Allaby and Pier Giorgio Innocenti for their important contributions to the LEPC during their terms as Division Leaders of PPE and ECP, respectively. They will be replaced from the next meeting by the new Division Leaders, Giorgio Goggi and Michal Turala.
 

10. Next LEPC meeting
 

The LEPC meeting on 31 January - 1 February has been cancelled. The dates of the next meeting of the LEPC are therefore
Tuesday and Wednesday, 6 - 7 June
 

                                                                                                                            J. Kirkby